# Gaia Data Release 3: spectroscopic binary-star orbital solutions

### v1

@@ 8.2.3 Problematic cases: fake SB1 and spectrocentre
We present here the illustrative case of HD 223323. This star ( = 6.968) is a well-known binary object of type SB2 that has been well studied from the ground; it has a period of 1175.1 d (Griffin 2007). The two components are rather similar, and only a merged spectral type in the range F2-4 IV-V is reported (see references in Griffin 2007). An accurate orbital solution has been published by the latter author, and the elements of interest for the present discussion are given in Table 7. Figure 10 illustrates the fitted orbital model published by Griffin (2007). The ground observations at phase 0.98 (almost at larger separation, see figure 5 of Griffin 2007) suggest that the lines of the two stars are almost deblended.
@@
| (km s-1)∗*∗∗*∗The systemic velocity is expressed in the Griffin RV reference system (see also Sect. 9.4.1) | –9.560.06 | –9.940.10 |
@@
The star has been observed with Gaia and is present in the DR3 catalogue under the name Gaia DR3 2852594583674129152. A total of 24 measurements () were collected by the RVS over a span of time of d. Two of the transits exhibit a double-line spectrum, and RVs have been derived as explained in Damerdji et al. (2024). For the remaining 22 transits, the star has been measured as a single-line object. Owing to the small percentage of time when the star appeared as double-lined, the pipeline decided to discard the two corresponding transits (see the treatment described by Damerdji et al. 2024). Therefore, the object entered the SB1 chain with data points. An SB1 fitted model was presented in the catalogue with a period 977.9 104.3 d. This value is in principle only indicative, since it is larger than the total span of time by a factor 1.11. In any case, the Gaia period is in good agreement with the precise Griffin period since it is only at 1.89 . The value of the period will automatically improve in the coming releases due to the increase of the . The other parameters of interest derived by Gaia are available in Table 7. The measured Gaia RVs are plotted in Fig. 10. The red dots represent the RVs derived by the single-line channel (they are put in the Griffin reference system by adding 1 km s-1, see Sect. 9.4.1 below). The other data points represent the primary (in blue) and the secondary (in cyan) RVs as derived in CU6 (Damerdji et al. 2024). They were also corrected to the Griffin reference zero point. It is remarkable that the Gaia solution yields a very small value for the semi-amplitude. The interpretation is quite straightforward. Except for the phases near the extrema in RVs, the resolution of the RVS and the predominance of the broad Calcium lines of the triplet makes that the lines never deblend at the other phases. After the rejection of the two double-line transits, the star becomes a fake SB1 which is actually a never-deblending SB2. There is a tendency for the position of the blended lines to moderately follow the behaviour of the primary variation. According to (Griffin 2007), the lines of the secondary are slightly broader and the lines of the primary slightly deeper. This thus implies that the primary lines are moderately dominating the blend. The secondary line depths are roughly 90 % of those of the primary one (we will call this ratio B = 0.9). The effect of the width of the lines of the secondary is attenuated since the lines are rather strongly blended. Tatum (1968) presented a computation of the position of the minimum of the blend of two Gaussian profiles (which is unique in the case of a strong blend) as a function of the widths and of the depths of the two profiles. Following their work, using the same notation as in their Eq. 5, and deciding to neglect the difference in broadness of the lines, we derived a position of the minimum which is given by
@@
It is interesting to notice that, in the case of HD 223323, the most affected parameter is essentially the semi-amplitude. On the opposite, it is clear from Table 7 that the other parameters are not strongly perturbed. The eccentricity in particular is well determined, as well as the argument of periastron. This is a consequence of the slight domination of the primary on the secondary effect.
@@ 8.2.4 Problematic cases: possible spuriously variable RVs
Some stars are associated with an orbital SB1-type solution that looks very convincing, and is identified as good according to the various quality tests. These stars thus enter the orbital solution catalogue. However, ground-based observations do not confirm these orbital variations. An illustrative case is the one of Gaia DR3 702159502868472832. Its RV time-series suggests a significant period of 1.0844000.000058 d with a probability of 0.997 and a resulting folded RV curve that is given in Fig. 11. With a significance of 15.41 and an , nothing allows us to doubt of the nature of this object. However, besides the Gaia RV curve and the fit, we plot also in the figure the few folded RVs that can be found in the literature from the APOGEE survey DR17 (Abdurro’uf et al. 2022). The two data sets are clearly not compatible. This suggests that some of the spectroscopic data from Gaia could not only present spurious RVs that are variable, but also that a period can be found from these measurements. This also suggests that the effect could be instrumental. The origin is currently not securely identified. A small number of objects enter this category, but the related sparsity could be due to the small percentage of objects having ground-based observations. This is supported by the validation performed in Sect. 9.4 below. Therefore, it is primordial to discover the origin in order to improve the robustness and purity of the future Data Release catalogues. We listed a few such doubtful cases that were detected here: they are Gaia DR3 66572718940795136, Gaia DR3 286652890498864768, Gaia DR3 487550398091020160, Gaia DR3 164423034699251968, Gaia DR3 577600125523574400, and the illustrative case Gaia DR3 702159502868472832. All of them have been compared to APOGEE DR17 data.
Another similar case has been detected by Bashi et al. (2022) but this time issued from comparison with the LAMOST RV survey (Cui et al. 2012, 999https://dr6.lamost.org). The object is Gaia DR3 3376949338201658112 and a significant periodicity has here also been identified ( d) and the corresponding RV curve is shown in Fig. 12. Again the RV variations are convincing as well as the fit. The object is one of the very few ones that have been observed at several epochs in the framework of the LAMOST survey. All the corresponding RVs are contained between the two green lines, further underlining the incompatibility of the ground-based data with the Gaia ones.
The fact that two ground-based surveys disagree with the results from Gaia underlines the fact that some instrumental effects are most probably present in the Gaia data. It is particularly intriguing that the RV time-series exhibit what could appear as a convincing periodic variability characterised by timescale not known to be problematic. Clearly, the problem is linked to the production of the RV time-series rather than to the behaviour of the algorithm of orbital parameter determination, except perhaps for the period.
@@ 9 Validation of the results
The analysis of the huge data set generated by Gaia and the establishment of the related catalogue necessitates some consolidation mechanism. This mechanism can be made of several approaches: they are described in the following subsections. The aim is evidently to test and fine-tune the pipeline itself in order to address its efficiency but also to discern the unavoidable biases. The understanding of the pipeline behaviour is directly related to the quality of the SB-subcatalogue and looking at its characteristics from all angles is a necessity. This validation process can be made on the basis of the internal consistency of the solutions but also on the basis of a comparison between the parameters delivered from the Gaia data with those obtained from the literature. This last approach is limited to the dramatically smaller amount of objects that have been the subject of a previous study. The validation paves the way for the determination of the global selection function.
@@ 9.1 Simulations
As a very first step, it is necessary to study the ability of the pipeline to recover the period. The study described here is based on the generation of synthetic RV curves of eccentric systems (with a random ), and of circular systems. Both sets of RV curves are generated with a random phase. We considered a number of data points of 10, 17, 30 and 80. The times of observations are mimicking the Gaia time series and are drawn in the same manner as in Sect. 4.4. The adopted signal-to-noise ratio is considered for values 5, 10 and 50. Fig. 13 reveals the recovery rate (in percent) for various true periods. The period is assumed to be recovered if its value does not differ from the input period by more than .
For equal to or larger than 10, the recovery rate is mostly better than 80 percent except when the number of data points is too low, see e.g. . This suggests that a too low number is a problem for the algorithm. It is also clear that an of 5 is equally detrimental. In Table 11, we give the corresponding value for the equivalent Significance given by
@@ 9.2.1 SB1 and SB1C classes
The analysis presented here below is essentially based on the output of the spectroscopic pipeline after the internal filtering. The SB1 data set contains 183434 objects. The basic reason for this choice is that the principal motivation is to first scrutinise the behaviour of the pipeline. The post-filtering is based on information from other channels and could thus be more complex to interpret, although it is necessary in order to constitute the DR3 SB-subcatalogue. In any case, for DR3, the differences between the analysed data set and the data set after post-filtering that contains 181327 objects, are almost negligible from a statistical point of view. The resulting data coming out of the pipeline were systematically inspected in order to verify the conformity to the expectation and the expected statistical distribution of the fitted parameters. Basically we are mainly treating the SB1 class excluding the SB1 solutions that were combined downstream with the solutions from other channels. The SB1C class contains too few objects to allow us to perform a fully sound statistical analysis. However, these solutions correspond to a very simple model that should not bear any particular problem.
@@
The distribution of the number of good transits (i.e. data points) for the SB1-type solutions processed by the pipeline is illustrated in Fig. 14. The objects having less than 10 transits were not further considered as explained above. This cut-off below 10 transits should not discard too many objects. Typically, we have slightly over 20 data points in our various time-series. The largest value is 228 (not shown here) but these objects are exceptional and are corresponding to the EPSL observations. As a first test, we examine the distribution of the magnitude of all solutions in the SB1 class. Figure 15 exhibits the histogram of for all the 183434 objects. We immediately notice that no object is present between and 5.5. This is an artefact of the selection based on the rv_renormalized_gof at the level of the input data. This quantity is defined when and thus below this threshold, the value for these objects contains NaN even if these bright objects should be kept. This is a weakness of DR3 and it will be corrected in future releases. On the other hand, it is also clear that signs of incompleteness are starting to appear above magnitude 11.5. A minimum of objects are lost between magnitude 11.75 and 12.0. This is most probably an effect of the increased importance that noise could have for the fainter objects, thus leading to badly determined solutions. However, an incompleteness also appears in the global RV output of the CU6 pipeline (see fig. 4 of Katz et al. 2023) for objects fainter than (corresponding for a solar-like star to ). Therefore, the difficulties to measure RVs on noisy spectra, the selection of variable objects in noisy time-series and the subsequent fit of orbital solutions conspire to induce this incompleteness.
@@
The observed distribution of the quality-of-fitting statistic deduced from the chi-square, the goodness of fit , is given for the filtered data set of solutions (see Fig. 16). The recommended and applied cut-off at is clearly visible. Otherwise, the original decay of the histogram is going up to values of several tens. Large values of the statistic are to be associated with bad, unacceptable solutions. This is certainly so for larger than 5. The cases between 3 and 5 could be further discussed. Conservatively, we only kept values below 3. This induces the rejection of some acceptable solutions, and consequently of the related objects. Along with the observed distribution, we also plot a Gaussian around the mean zero and with a dispersion (in magenta). This curve represents the expected behaviour of under the null-hypothesis of a perfect Keplerian model with additive Gaussian noise (corresponding to the uncertainties associated to the individual RVs). Another Gaussian curve with a (in blue) is also shown. This latter better follows the histogram on its left tail. The side of the negative values of corresponds to solutions for which the realisation of the noise conspires to produce better fitting. On the other hand, the right tail is associated to the opposite effect where the solutions are noisier although still acceptable. It is interesting to notice that, in this region, the histogram reveals a remarkably too large number of objects. This is true even comparing to the blue Gaussian curve; an excess is thus present. They most probably correspond to solutions for objects following a Keplerian orbit (thus generating a rather good fit) but affected by additional variations (pulsations, activity, spots, …), not due to the observational noise, but at a sufficiently low level such that the dominating effect is still the orbital variation. Among these objects, multiple systems and intrinsic variables could be the main contributors to this effect. A more sophisticated study, beyond the scope of the present paper, is necessary. However, one should not be misled by the normalisation of the comparison Gaussian curves. Actually, if we normalised with respect to the area, there would appear a deficit (roughly 25%) of objects with solutions characterised by an broadly located around zero (essentially between -1. and +1.). These solutions are displaced towards the right tail. The effect is present both on the individual (with a selected value of , the number of degrees of freedom) and on the global statistics. The transformation of the values into the global statistic is not the cause of the deviation.
@@
Another interesting plot is the distribution of versus the period. It is given in Fig. 19. Since is a cyclic variable of modulo (see Sect. 3), an infinity of possible values are available. The data points being located around , the value presenting the best precision and a minimum correlation is necessarily situated between and . It is clear that the graph is essentially as expected with a rather homogeneous density of data points slowly decreasing towards the largest periods. For periods around 900 to 1100 days, there is a slight tendency to have a larger density of solutions near the values . A part of the solutions located there correspond to low significance and rather large eccentricities. This is dominantly an artefact of the decision we took to enforce an SB1 solution in this region of periods against the possibility to consider trends. Indeed, let us consider a time-series of a noisy linear trend. The enforced SB1 solution will deliver a period roughly about (only one ‘cycle’ but this depends on the exact sampling). The trend folded in phase will very often mimic an eccentric curve with a jump between the last data point of the time-series and the first one. Therefore, this jump, after folding, will be associated to the passage at periastron and, consequently, the returned will be attributed a value around . This conclusion is further supported by the 2D distributions of solutions (i.e. versus and ) shown in Fig. 20. One clearly sees that the excess of solutions is effectively due to the range of periods enforced to an SB1 solution (prominent zone in yellow in the bottom right panel, not present on the upper left and right ones). It is also clear that the effect is absent for quasi-circular orbits (which are fully correct) and is associated to larger eccentricities. In addition, one clearly sees that the solutions at the basis of the excess are confined to values of around and (very same zone in yellow). This indicates that the trends enforced into an SB1 solution generate a RV curve that is predominantly of the saw-tooth type, further confirming our conclusions. The excess partly contributes to the excess at specific noticed in Fig. 18. This problem will be corrected for DR4; it will also occur at another period since will be larger. Some other effects might also be present. Note that not all the solutions in this region are problematic.
@@
The critical point when fitting the orbital SB1 models is the determination of the period (see above). Fig. 25 displays the histogram of the period (in log). Several features are visible. The log-distribution is rather flat, with a marked excess at very large periods. This excess can be due to the fact (see above) that the periods of the same order of magnitude as the span of time of the mission are rather indicative and certainly not a perfect estimator of the true period. This uncertainty implies the concentration of several solutions near these long periods. This excess is very sensitive to the threshold separating SB1 and TrendSB1 solutions. Another explanation is the fact that the orbital solutions are enforced in this region. On the other hand, the deficit of solutions around the period of 182.6 days is due to the yearly Earth orbit that the Gaia satellite follows by staying at L2. The satellite avoiding the direction of the Sun, this produces regular gaps with a timescale of half a year. These gaps presented by the sampling are associated to large gaps in phase leading to the removal of the solutions because they correspond to badly determined periods and orbits. Fig. 27 is showing a diagram of the value of the largest gap in the solution as a function of the period. Exceptionally, here we use the data set of 367328 objects constituting the set of SB1 solutions before the application of the internal filtering. It is clear from this figure that some of the solutions at periods near 182.6 d generate phase diagrams that are associated to the presence of vast gaps. This situation seems also to be the case, but to a lesser extent, for the period around one year. The phenomenon is perhaps also present for periods of 91.3 d. A detailed look at this Fig. 27 suggests that a conservative choice for the threshold on the gap of 0.3 removed all the potentially spurious or at least unsafe solutions. The decrease of density of solutions for lower periods, as seen in Fig. 25, is an artefact of the application of an ad hoc selection criterion (see the paragraph on the internal filtering of the period in Sect. 7.1) that prevents the proliferation of the small periodicities. This effect should be further investigated for the next Data Releases.
@@
Figure 28 presents the histogram of the spectroscopic significance . The significance is distributed as a decreasing power law vanishing at 800. The maximum accepted value for being 250 km s-1, the maximum significance value corresponds to a typical uncertainty of about 0.3 km s-1 which is coherent with expectation. The distribution shows no inflection locus that could help to set a selection threshold. In this context, we decided to introduce a threshold at 5 but not more restrictive ones. Below a value of 5, the corresponding solutions are clearly not trustworthy. We leave the possibility to the catalogue users to define a threshold that could be appropriate for their own work.
Fig. 29 exhibits the distribution of the standard deviations on the eccentricity. The typical uncertainty on the eccentricity is of 0.03-0.10 . It is necessary to draw the attention to the fact that with such an uncertainty a derived eccentricity of 0.3-0.4 is still compatible with a circular orbit (this is within three keeping in mind that the distribution of eccentricities is not Gaussian). RVs are well-known to be a poor source of determination of eccentricities in contrast to the astrometric orbital solutions. Errors larger than 0.35-0.4 (there are a few) could be associated to insecure adjustments and we encourage the catalogue user to be careful with these questionable solutions.
@@
Finally, Fig. 31 represents the plot of the derived SB1 mass function as a function of the adopted period. One clearly sees that the majority of the data points are between 0.001 and 0.5 M☉. Objects outside this range are very few, in particular on the side of the high mass function. From the statistical point of view, the diagram is rather close to the expected one. Data points at an unexpected location could be due to problems in the derivation of the period or of the eccentricity. In particular, one should be very careful while searching for abnormal mass function values or massive black holes. In addition, the overdensities tracing vertical lines at days and at one year (much broader) are due to objects with a varying apparent position inducing systematic effects on the RVs associated to the variation of the scan-angle. These are objects presenting small amplitudes at these particular periodicities (bad astrometry inducing a wrong wavelength calibration). These objects could easily be filtered out but they were not before the application of the combiner with the astrometric channel in order not to miss interesting objects. The effect in astrometry is described in Lindegren (2019). A more detailed analysis of the scan-angle dependent signals is given in Holl et al. (2023a).

### v2

@@ 8.2.3 Problematic cases: fake SB1 and spectrocentre
We present here the illustrative case of HD 223323. This star ( = 6.968) is a well-known binary object of type SB2 that has been well studied from the ground; it has a period of 1175.1 d (Griffin 2007). The two components are rather similar, and only a merged spectral type in the range F2-4 IV-V is reported (see references in Griffin 2007). An accurate orbital solution has been published by the latter author, and the elements of interest for the present discussion are given in Table 7. Figure 10 illustrates the fitted orbital model published by Griffin (2007). The ground observations at phase 0.98 (almost at larger separation; see Fig. 5 of Griffin 2007) suggest that the lines of the two stars are almost deblended.
@@
| (km s-1)∗*∗∗*∗The systemic velocity is expressed in the Griffin RV reference system (see also Sect. 9.4.1). | –9.560.06 | –9.940.10 |
@@
The star has been observed with Gaia and is present in the DR3 catalogue under the name Gaia DR3 2852594583674129152. A total of 24 measurements () were collected by the RVS over a span of time of d. Two of the transits exhibit a double-line spectrum, and RVs have been derived as explained in Damerdji et al. (2024). For the remaining 22 transits, the star has been measured as a single-line object. Owing to the small percentage of time when the star appeared as double-lined, the pipeline decided to discard the two corresponding transits (see the treatment described by Damerdji et al. 2024). Therefore, the object entered the SB1 chain with data points. An SB1 fitted model was presented in the catalogue with a period 977.9 104.3 d. This value is in principle only indicative, since it is larger than the total span of time by a factor 1.11. In any case, the Gaia period is in good agreement with the precise Griffin period since it is only at 1.89 . The value of the period will automatically improve in the coming releases due to the increase of the . The other parameters of interest derived by Gaia are available in Table 7. The measured Gaia RVs are plotted in Fig. 10. The red dots represent the RVs derived by the single-line channel (they are put in the Griffin reference system by adding 1 km s-1; see Sect. 9.4.1 below). The other data points represent the primary (in blue) and the secondary (in cyan) RVs as derived in CU6 (Damerdji et al. 2024). They were also corrected to the Griffin reference zero point. It is remarkable that the Gaia solution yields a very small value for the semi-amplitude. The interpretation is quite straightforward. Except for the phases near the extrema in RVs, the resolution of the RVS and the predominance of the broad Calcium lines of the triplet makes that the lines never deblend at the other phases. After the rejection of the two double-line transits, the star becomes a fake SB1 which is actually a never-deblending SB2. There is a tendency for the position of the blended lines to moderately follow the behaviour of the primary variation. According to Griffin (2007), the lines of the secondary are slightly broader and the lines of the primary slightly deeper. This thus implies that the primary lines are moderately dominating the blend. The secondary line depths are roughly 90 % of those of the primary one (we will call this ratio B = 0.9). The effect of the width of the lines of the secondary is attenuated since the lines are rather strongly blended. Tatum (1968) presented a computation of the position of the minimum of the blend of two Gaussian profiles (which is unique in the case of a strong blend) as a function of the widths and of the depths of the two profiles. Following their work, using the same notation as in their Eq. 5, and deciding to neglect the difference in broadness of the lines, we derived a position of the minimum which is given by
@@
It is interesting to notice that in the case of HD 223323, the most affected parameter is essentially the semi-amplitude. On the contrary, it is clear from Table 7 that the other parameters are not strongly perturbed. The eccentricity in particular is well determined, as well as the argument of periastron. This is a consequence of the slight domination of the primary on the secondary effect.
@@ 8.2.4 Problematic cases: possible spuriously variable RVs
Some stars are associated with an orbital SB1-type solution that looks very convincing, and is identified as good according to the various quality tests. These stars thus enter the orbital solution catalogue. However, ground-based observations do not confirm these orbital variations. An illustrative case is the one of Gaia DR3 702159502868472832. Its RV time series suggests a significant period of 1.0844000.000058 d with a probability of 0.997 and a resulting folded RV curve that is given in Fig. 11. With a significance of 15.41 and an , there is no solid reason to doubt the nature of this object. However, besides the Gaia RV curve and the fit, we plot also in the figure the few folded RVs that can be found in the literature from the APOGEE survey DR17 (Abdurro’uf et al. 2022). The two data sets are clearly not compatible. This suggests that some of the spectroscopic data from Gaia could not only present spurious RVs that are variable, but also that a period can be found from these measurements. This also suggests that the effect could be instrumental. The origin is currently not securely identified. A small number of objects enter this category, but the related sparsity could be due to the small percentage of objects having ground-based observations. This is supported by the validation performed in Sect. 9.4 below. Therefore, it is primordial to discover the origin in order to improve the robustness and purity of the future Data Release catalogues. We listed a few such doubtful cases that were detected here: they are Gaia DR3 66572718940795136, Gaia DR3 286652890498864768, Gaia DR3 487550398091020160, Gaia DR3 164423034699251968, Gaia DR3 577600125523574400, and the illustrative case Gaia DR3 702159502868472832. All of them have been compared to APOGEE DR17 data.
Another similar case has been detected by Bashi et al. (2022) but this time issued from comparison with the LAMOST RV survey (Cui et al. 2012, 999https://dr6.lamost.org). The object is Gaia DR3 3376949338201658112 and a significant periodicity has here also been identified ( d) and the corresponding RV curve is shown in Fig. 12. Again the RV variations are convincing as well as the fit. The object is one of very few that have been observed at several epochs in the framework of the LAMOST survey. All the corresponding RVs are contained between the two green lines, further underlining the incompatibility of the ground-based data with the Gaia ones.
The fact that two ground-based surveys disagree with the results from Gaia underlines the fact that some instrumental effects are most probably present in the Gaia data. It is particularly intriguing that the RV time series exhibit what could appear as a convincing periodic variability characterised by timescale not known to be problematic. Clearly, the problem is linked to the production of the RV time series rather than to the behaviour of the algorithm of orbital parameter determination, except perhaps for the period.
@@ 9 Validation of the results
The analysis of the huge data set generated by Gaia and the establishment of the related catalogue necessitates some consolidation mechanism. This mechanism can be made of several approaches: they are described in the following subsections. The aim is evidently to test and fine-tune the pipeline itself in order to address its efficiency but also to discern the unavoidable biases. The understanding of the pipeline behaviour is directly related to the quality of the SB subcatalogue and looking at its characteristics from all angles is a necessity. This validation process can be made on the basis of the internal consistency of the solutions but also on the basis of a comparison between the parameters delivered from the Gaia data with those obtained from the literature. This last approach is limited to the dramatically smaller amount of objects that have been the subject of a previous study. The validation paves the way for the determination of the global selection function.
@@ 9.1 Simulations
As a very first step, it is necessary to study the ability of the pipeline to recover the period. The study described here is based on the generation of synthetic RV curves of eccentric systems (with a random ), and of circular systems. Both sets of RV curves are generated with a random phase. We considered a number of data points of 10, 17, 30, and 80. The times of observations are mimicking the Gaia time series and are drawn in the same manner as in Sect. 4.4. The adopted S/N ratio is considered for values 5, 10, and 50. Figure 13 reveals the recovery rate (in percent) for various true periods. The period is assumed to be recovered if its value does not differ from the input period by more than .
For equal to or larger than 10, the recovery rate is mostly better than 80 percent except when the number of data points is too low (see e.g. . This suggests that a too low number is a problem for the algorithm. It is also clear that an of 5 is equally detrimental. In Table 11, we give the corresponding value for the equivalent significance given by
@@ 9.2.1 SB1 and SB1C classes
The analysis presented here below is essentially based on the output of the spectroscopic pipeline after the internal filtering. The SB1 data set contains 183 434 objects. The basic reason for this choice is that the principal motivation is to first scrutinise the behaviour of the pipeline. The post-filtering is based on information from other channels and could thus be more complex to interpret, although it is necessary in order to constitute the DR3 SB subcatalogue. In any case, for DR3, the differences between the analysed data set and the data set after post-filtering that contains 181 327 objects, are almost negligible from a statistical point of view. The resulting data coming out of the pipeline were systematically inspected in order to verify the conformity to the expectation and the expected statistical distribution of the fitted parameters. Basically we are mainly treating the SB1 class excluding the SB1 solutions that were combined downstream with the solutions from other channels. The SB1C class contains too few objects to allow us to perform a fully sound statistical analysis. However, these solutions correspond to a very simple model that should not bear any particular problem.
@@
The distribution of the number of good transits (i.e. data points) for the SB1-type solutions processed by the pipeline is illustrated in Fig. 14. The objects having less than 10 transits were not further considered as explained above. This cut-off below 10 transits should not discard too many objects. Typically, we have slightly over 20 data points in our various time series. The largest value is 228 (not shown here) but these objects are exceptional and are corresponding to the EPSL observations. As a first test, we examine the distribution of the magnitude of all solutions in the SB1 class. Figure 15 exhibits the histogram of for all the 183 434 objects. We immediately notice that no object is present between and 5.5. This is an artefact of the selection based on the rv_renormalised_gof at the level of the input data. This quantity is defined when and thus below this threshold, the value for these objects contains NaN even if these bright objects should be kept. This is a weakness of DR3 and it will be corrected in future releases. On the other hand, it is also clear that signs of incompleteness are starting to appear above magnitude 11.5. A minimum of objects are lost between magnitude 11.75 and 12.0. This is most probably an effect of the increased importance that noise could have for the fainter objects, thus leading to badly determined solutions. However, an incompleteness also appears in the global RV output of the CU6 pipeline (see fig. 4 of Katz et al. 2023) for objects fainter than (corresponding for a solar-like star to ). Therefore, the difficulties to measure RVs on noisy spectra, the selection of variable objects in noisy time series and the subsequent fit of orbital solutions conspire to induce this incompleteness.
@@
The observed distribution of the quality-of-fitting statistic deduced from the chi-square, the goodness of fit , is given for the filtered data set of solutions (see Fig. 16). The recommended and applied cut-off at is clearly visible. Otherwise, the original decay of the histogram is going up to values of several tens. Large values of the statistic are to be associated with bad, unacceptable solutions. This is certainly so for larger than 5. The cases between 3 and 5 could be further discussed. Conservatively, we only kept values below 3. This induces the rejection of some acceptable solutions, and consequently of the related objects. Along with the observed distribution, we also plot a Gaussian around the mean zero and with a dispersion (in magenta). This curve represents the expected behaviour of under the null-hypothesis of a perfect Keplerian model with additive Gaussian noise (corresponding to the uncertainties associated to the individual RVs). Another Gaussian curve with a (in blue) is also shown. This latter better follows the histogram on its left tail. The side of the negative values of corresponds to solutions for which the realisation of the noise conspires to produce better fitting. On the other hand, the right tail is associated to the opposite effect where the solutions are noisier although still acceptable. It is interesting to notice that, in this region, the histogram reveals a remarkably too large number of objects. This is true even comparing to the blue Gaussian curve; an excess is thus present. They most probably correspond to solutions for objects following a Keplerian orbit (thus generating a rather good fit) but affected by additional variations (pulsations, activity, spots, …), not due to the observational noise, but at a sufficiently low level such that the dominating effect is still the orbital variation. Among these objects, multiple systems and intrinsic variables could be the main contributors to this effect. A more sophisticated study, beyond the scope of the present paper, is necessary. However, we should not be misled by the normalisation of the comparison Gaussian curves. Actually, if we normalised with respect to the area, there would appear a deficit (roughly 25%) of objects with solutions characterised by an broadly located around zero (essentially between –1. and +1.). These solutions are displaced towards the right tail. The effect is present both on the individual (with a selected value of , the number of degrees of freedom) and on the global statistics. The transformation of the values into the global statistic is not the cause of the deviation.
@@
Another interesting plot is the distribution of versus the period. It is given in Fig. 19. Since is a cyclic variable of modulo (see Sect. 3), an infinity of possible values are available. The data points being located around , the value presenting the best precision and a minimum correlation is necessarily situated between and . It is clear that the graph is essentially as expected with a rather homogeneous density of data points slowly decreasing towards the largest periods. For periods around 900 to 1100 days, there is a slight tendency to have a larger density of solutions near the values . A part of the solutions located there correspond to low significance and rather large eccentricities. This is dominantly an artefact of the decision we took to enforce an SB1 solution in this region of periods against the possibility to consider trends. Indeed, let us consider a time series of a noisy linear trend. The enforced SB1 solution will deliver a period roughly about (only one ‘cycle’ but this depends on the exact sampling). The trend folded in phase will very often mimic an eccentric curve with a jump between the last data point of the time series and the first one. Therefore, this jump, after folding, will be associated to the passage at periastron and, consequently, the returned will be attributed a value around . This conclusion is further supported by the 2D distributions of solutions (i.e. versus and ) shown in Fig. 20. We can clearly see that the excess of solutions is effectively due to the range of periods enforced to an SB1 solution (prominent zone in yellow in the bottom right panel, not present on the upper left and right ones). It is also clear that the effect is absent for quasi-circular orbits (which are fully correct) and is associated to larger eccentricities. In addition, we can clearly see that the solutions at the basis of the excess are confined to values of around and (very same zone in yellow). This indicates that the trends enforced into an SB1 solution generate a RV curve that is predominantly of the saw-tooth type, further confirming our conclusions. The excess partly contributes to the excess at specific noticed in Fig. 18. This problem will be corrected for DR4; it will also occur at another period since will be larger. Some other effects might also be present. Note that not all the solutions in this region are problematic.
@@
The critical point when fitting the orbital SB1 models is the determination of the period (see above). Figure 25 displays the histogram of the period (in log). Several features are visible. The log-distribution is rather flat, with a marked excess at very large periods. This excess can be due to the fact (see above) that the periods of the same order of magnitude as the span of time of the mission are rather indicative and certainly not a perfect estimator of the true period. This uncertainty implies the concentration of several solutions near these long periods. This excess is very sensitive to the threshold separating SB1 and TrendSB1 solutions. Another explanation is the fact that the orbital solutions are enforced in this region. On the other hand, the deficit of solutions around the period of 182.6 days is due to the yearly Earth orbit that the Gaia satellite follows by staying at L2. The satellite’s avoidance of the direction of the Sun produces regular gaps with a timescale of half a year. These gaps presented by the sampling are associated to large gaps in phase leading to the removal of the solutions because they correspond to badly determined periods and orbits. Figure 27 is showing a diagram of the value of the largest gap in the solution as a function of the period. Exceptionally, here we use the data set of 367 328 objects constituting the set of SB1 solutions before the application of the internal filtering. It is clear from this figure that some of the solutions at periods near 182.6 d generate phase diagrams that are associated to the presence of vast gaps. This situation seems also to be the case, but to a lesser extent, for the period around one year. The phenomenon is perhaps also present for periods of 91.3 d. A detailed look at this Fig. 27 suggests that a conservative choice for the threshold on the gap of 0.3 removed all the potentially spurious or at least unsafe solutions. The decrease of density of solutions for lower periods, as seen in Fig. 25, is an artefact of the application of an ad hoc selection criterion (see the paragraph on the internal filtering of the period in Sect. 7.1) that prevents the proliferation of small periodicities. This effect should be further investigated in the next data releases.
@@
Figure 28 presents the histogram of the spectroscopic significance . The significance is distributed as a decreasing power law vanishing at 800. The maximum accepted value for being 250 km s-1, the maximum significance value corresponds to a typical uncertainty of about 0.3 km s-1 which is coherent with expectation. The distribution shows no inflection locus that could be useful in setting a selection threshold. In this context, we decided to introduce a threshold at 5 but not more restrictive ones. Below a value of 5, the corresponding solutions are clearly not trustworthy. We leave the possibility to the catalogue users to define a threshold that could be appropriate for their own work.
Figure 29 exhibits the distribution of the standard deviations on the eccentricity. The typical uncertainty on the eccentricity is of 0.03-0.10 . It is necessary to draw the attention to the fact that with such an uncertainty a derived eccentricity of 0.3-0.4 is still compatible with a circular orbit (this is within three keeping in mind that the distribution of eccentricities is not Gaussian). RVs are well-known to be a poor source of determination of eccentricities in contrast to the astrometric orbital solutions. Errors larger than 0.35-0.4 (there are a few) could be associated to insecure adjustments and we encourage the catalogue user to be careful with these questionable solutions.
@@
Finally, Fig. 31 represents the plot of the derived SB1 mass function as a function of the adopted period. We can clearly see that the majority of the data points are between 0.001 and 0.5 M☉. Objects outside this range are very few, in particular on the side of the high mass function. From the statistical point of view, the diagram is rather close to the expected one. Data points at an unexpected location could be due to problems in the derivation of the period or of the eccentricity. In particular, we should be very careful while searching for abnormal mass function values or massive black holes. In addition, the overdensities tracing vertical lines at days and at one year (much broader) are due to objects with a varying apparent position inducing systematic effects on the RVs associated to the variation of the scan-angle. These are objects presenting small amplitudes at these particular periodicities (bad astrometry inducing a wrong wavelength calibration). These objects could easily be filtered out but they were not before the application of the combiner with the astrometric channel in order not to miss interesting objects. The effect in astrometry is described in Lindegren (2019). A more detailed analysis of the scan-angle dependent signals is given in Holl et al. (2023a).